# Are language models rational? The case of coherence norms and belief revision

### v2

@@ 3 Coherence norms for language models
Belief in humans is subject to the norms of rationality. We are not always rational, but we ought to be, which is to say: we are under the norms of rationality, even if we don’t always live up to those requirements. Rationality itself is complex, and the requirement to be rational can be analyzed as a series of distinct norms. One core part of these norms are coherence norms, which concern how the internal states of a system ought to relate to each other. Other norms of rationality concern the proper response to reasons or to evidence. How the different parts of rationality relate to each other is itself controversial, but we will simply focus on coherence norms in the following. [^4] See Kolodny (2005), Kiesewetter (2017), Worsnip (2021) and many others for the debate about whether coherence norms are derivative on norms tied to responding to reasons or whether they are an independent group of norms of rationality.
@@
Coherence norms come in different kinds, most notably as ones concerning full belief and ones concerning degrees of, or strength of belief. Most notably among the former are norms concerning logical coherence and among the latter are norms about probabilistic coherence. Logical coherence norms concern the relation of our beliefs to each other with respect to logical implication and consistency. To give a simple example, a logical coherence norm might demand that one ought not to believe a particular proposition $p$ as well as its negation $\neg p$. Probabilistic coherence norms in general also come in different kinds. One kind concerns belief about probabilities itself. For example, a norm might demand that one ought not to believe that the probability that it is going to rain is 90%, and also believe that the probability that it is not going to rain is 60%. But besides beliefs explicitly about probabilities, there are also coherence norms tied to subjective probabilities or credences. Credences correspond to how strongly one believes a particular proposition. So, to believe that Paris is in France with credence 99% is to believe it very strongly, but it is not itself a belief about probabilities. Therefore to believe that Paris is in France with 99% credence is not the same as to believe that the chance that Paris is in France is 99%. Credences should be coherent, and the view that they should satisfy the axioms of probability theory is generally called probabilism. For example, probabilism holds that if your credence that $p$ is $x$, then your credence that $\neg p$ ought to be $(1-x)$, for $x\in[0,1]$. We will not discuss coherence norms tied to beliefs explicitly about probabilities in this paper, but focus on logical coherence and the coherence of credences instead.
@@
First, it is unclear why logical coherence should be required of language models. After all, they are machine learning models and thus they are supposed to learn the training data. But the training data for language models is generally incoherent. Language models are trained on text from the internet and other sources, which is not a coherent unified dataset of text. Since the machine learning model is trained on incoherent text as its training data, why should we demand of it that it be coherent? In fact, coherence seems to go against what it is supposed to learn from the training data. Incoherent training data should lead to a model with incoherent beliefs if the model properly captures the training data. Thus to require the model to be coherent seems to be unjustified, and therefore the rational norm of coherence does not seem to apply to it. Unless the training data is coherent the model should not be required to be coherent either. This contrasts with why coherence norms apply to human beings. We learn from a coherent world: a world without any contradictions. Thus our internal states should be coherent, to properly reflect what our minds are trained on. Even if our evidence is conflicting, pointing partly in one direction and partly in another one, what the evidence points to — the world — is coherent. Thus the norm of logical coherence applies to us, since we aim to represent a coherent world.
@@ 3.1 Logical coherence norms
The solution to the first problem is similar to the argument given above in Section 2 for the coherence of beliefs in language models. It is true in general that a pre-trained language model should not be seen as being under the norm of logical coherence. Such a model is simply aiming to model the training data, and that data is not logically coherent. To require coherence of a pre-trained language model that is trained on incoherent data is to require something of it that goes against what it fundamentally aims to do: learn the training data. But things are again different for fine-tuned models, in particular ones that are fine-tuned for truthfulness. Properly fine-tuned models can acquire the function to produce truthful text, and if we assume that such models have internal representational states like beliefs and that those states are properly causally connected to the text they produce, then the aim of being truthful can support that logical coherence norms apply to the internal states of such models. Internal belief states that are incoherent will be contrary to the norm of truth which is part of the standard of correctness for belief itself. Thus inconsistent beliefs are in conflict with the standard of correctness of belief, and thus the coherence norm of consistency for beliefs applies. A pre-trained model does not need to be under such norms, but a model fine-tuned for truthfulness is.
@@ 3.2 Defining credence for language models
which are assigned by the language model to the n-th token given that the previous tokens are $t_{1},...,t_{n-1}$, and turn that into an account of what the models credence in proposition $p$ is.
@@
We propose to solve this problem via the Minimal Assent Connection (MAC). The basic idea of MAC is simply to consider the probability of the model predicting the next token to be ‘yes’ when asked whether it is the case that $p$. This way next token probabilities can be directly tied to probabilities assigned to propositions. To put a label on it, we can say that the model minimally assents to $p$ just in case the probability it assigns to ‘yes’ being the next token after ‘Is it the case that $p$?’ is sufficiently high. This basic idea is promising, but to in order to turn it into a credence function that captures strength of belief more needs to be said and we will do this momentarily. But on the positive side, we can already note that this approach is completely general, since for any proposition $p$ we can consider the probabilities that the model assigns to the next token after the question whether it is the case that $p$. This probability comes from the internals of the language model as a machine learning model, and it is just the kind of probability we want to exploit in assigning credences to the model. This assignment thus applies to any proposition $p$ uniformly. Furthermore, the probability of a ‘yes’ answer to the question whether $p$ is naturally associated with the strength of believing $p$. After all, it is natural to hold that the probability assigned to ‘yes’ after the question whether $p$ is closely correlated with the strength of believing $p$: the stronger the confidence of the model in $p$, the higher the probability for a ‘yes’ answer, all things being equal. But still, it can’t quite be a full answer and we can’t make this connection as outlined. That is, we can’t fully endorse that we can define $cr(p)$, the models credence in $p$, as follows:
$cr(p)=P(\text{Yes }\mid\text{ Is it the case that }p?)$ (1)
@@
There are essentially two flaws with this. First, the model might assent to $p$ in ways other than with ‘yes’. It might also assent via other forms of affirmation, say ‘yeah’ or ‘indeed’. Using just (1) thus gives a credence that is too low. But we can simply fix this in one of two ways. First, we could force the model to answer either ‘yes’ or ‘no’, either by adding this requirement to the question prompt, or by constraining the model in other ways to do so (e.g. by constraining and re-normalizing the output distribution). However, adding such further constraints slightly changes the model or the text for which it predicts the next token. Making such a demand for either a ‘yes’ or a ‘no’ answer thus requires changes that make the method not completely general any more. It wouldn’t apply to any model any longer, only to ones properly modified to meet this constraint. But there is also a second way to solve this problem. We can leave the model and input alone, and instead simply sum over all affirmations, not just ‘yes’, but also ‘sure’, ‘yeah’, and so on, including multi-token sequences like ‘I couldn’t agree more.’ All such tokens or sequences of tokens we can label an assent sequence. If we call $AS$ the set of all such assent sequences, then we can better define the credence for $p$ by the model as the sum of all conditional probabilities of assenting. So, we can define the assent probability as:
@@
But we can fix the approach to overcome this difficulty. Strength of belief can be associated with how likely the model is to minimally assent when compared to how likely it is to minimally dissent, leaving aside other ways to respond to the question besides assenting or dissenting. So, we can simply consider the ration of assenting to either assenting or dissenting. To make this idea work, we also need to consider the set of all dissent sequences $DS$, which are the ways the model might minimally dissent: ‘no’, ‘never’, etc..
@@
So understood we should not expect that $as(p)$ and $ds(p)$ sum to 1, since the model might respond in other ways than assenting or dissenting. But we can simply re-normalize by considering only the ratio of probabilities given to affirmations to those that are given to either affirmations or dissent, leaving out all other options. And this ratio is the proper way to assess the strength of belief of a model in a proposition $p$: how likely it is to assent when we consider only assent or dissent, leaving out all other options. This then is the proposal for how to assign credences to a language model in a way that is derivative only on the model’s own assignment of probabilities to the next token:
@@
We have not given a detailed list of assent and dissent sequences, and doing so requires one to make some choices that can be subject to debate. One concern is tied to context sensitivity. It could be that what counts as assent or dissent for a particular issue $p$ depends itself on $p$, and thus it is not clear whether assent and dissent sequences can simply be listed for questions concerning any $p$. Although this is in principle possible, it does not seem to be a major concern. First, it is unclear what plausible cases there are for this, and second it merely requires a reformation of the proposal by relativizing assent and dissent sequences to a given $p$, and then consider the assent/dissent ratios concerning $p$ with with each those relativized to $p$ itself.
A more difficult case is presented by examples of sequences that concern the epistemic state of the model itself, as ‘It seems to me to be unlikely,’ ‘I am quite sure that $p$,’ or ‘it it is doubtful.’ Such sequences contain ‘epistemic markers’ or ‘epistemic strengtheners or weakeners’ that report on the epistemic state of the model. Strong epistemic markers indicate that the model is in a strong epistemic position to make an assessment, as with ‘I am sure that…,’ or ‘I am quite certain that..’. Weak epistemic markers indicate a weak epistemic state, as in ‘I am not sure, but it seems that…’ or ‘I doubt that…’. Should sequences with such epistemic markers, either strong or weak, count as assent or dissent sequences for our purposes here? Should next token probabilities allocated to such sequences count towards assent or dissent probabilities?
@@
Whether or not present day models’ credences match their epistemic markings is an empirical question. If they do they are epistemically better functioning than if they don’t. To determine whether there indeed is this correlation is subject to ongoing work. See also Zhou et al. (2024) for more on models reporting on their own epistemic state.
Furthermore, we would like to make clear that the assignment of credences to beliefs for models as defined above are for revealed belief, since they concern the model’s response to questions. This approach will not work for models that have developed the ability to lie and thus answer questions contrary to what they believe. How to make sense of lying in language models is not clear, since it is not clear how to make sense of language model belief besides revealed belief. To be clear, lying here is understood not simply as asserting something false, or even as disregarding the truth altogether, but as asserting something that is contrary to what one believes. To make sense of lying on this understanding of it requires that one can make sense of belief in language models independently of the model’s produced text. Whether this can be done is unclear, and our approach does not aim to do this, but focuses on revealed belief instead.
Finally, we would like to note that
@@ 3.3 The argument for probabilism
When focusing on synchronic coherence norms, this question quickly turns into the question whether the credence function should satisfying the axioms of probability theory and thus be a probability function. In the philosophical literature the view that a credence function should be a probability function is called probabilism. [^5] See Joyce (1998), Teitelbaum (2023). Assuming credences for language models as spelled out above, does probabilism apply to them as well? Are language models under the coherence norm that their credence function satisfies the axioms of probability theory? For example, are they under the norm that $cr(p)=1-cr(\neg p)$?
To answer this question we first need to again distinguish pre-trained language models from fine-tuned ones, in particular those fine-tuned for truthfulness. For a pre-trained model we should not in general expect the model’s credences to obey the probability calculus, nor that they are under any norm to do so. Since the model will generally be trained on incoherent text, we can not expect that $cr(p)$, the credence assigned to $p$, is exactly $1-cr(\neg p)$. If the training data is sufficiently incoherent, then it can well be that the credence for each of them is over 50%. Since the model so far is merely a pure language model whose internal states are correct insofar they contribute to the training task and learning the data, we should not expect there to be a norm on these probability assignments that correspond to probabilism.
However, things are again different for the model that is fine-tuned on truthfulness. A model that aims at true beliefs is subject to the well-known accuracy arguments for probabilism. See Joyce (1998), Teitelbaum (2023). The main idea of these arguments is as follow: if the goal is truth, then credences should be as close to the truth as possible: one’s credence in a true proposition should be close to 1, and one’s credence in a false proposition should be close to 0. However close one comes to this ideal can be seen as the accuracy of one’s credences. A standard measure of accuracy in this sense is the Brier score, basically the Euclidean distance between one’s credences in some propositions and their actual truth values (taken as 1 for true, and 0 for false) seen as points in a Euclidean space of the dimension that corresponds to the number of propositions, but without taking a square root. So, the Brier score for a credence function $cr$ from a finite set of propositions $p_{1},...p_{n}$ to [0,1] is defined as
@@
whereby $tv(p)=1$ if p is true and 0 if it is false. [^6] The Brier score is sometimes also given as the mean squared error, i.e. dividing the above by the dimension $n$. This difference does not matter for us here. The thought is that a credence function is better if its Brier score is lower, since one aims at the truth, and a lower Brier score of one’s credence function means that one’s credences are closer to how they should be. In Joyce (1998), James Joyce showed that if one’s credence function does not meet the axioms of probability theory, then there is a different credence function that has a lower Brier score no matter what the world is like and which meets the axioms of probability theory. However, if one’s credence function does meet the axioms of probability theory, then this is not always the case. This argument is taken to support probabilism: A credence function that isn’t a probability function can always be dominated by one that is. Thus such a credence function isn’t what it should be, given that one aims at the truth. And thus probabilism is a rational norm that applies to ones credences: one’s credence function ought to be a probability function. [^7] For further discussion of Joyce’s argument, see Teitelbaum (2023) volume 2, pages 338-71. This argument applies to the credence functions of language models just as it applies to the credence functions of human beings. As long as we aim at the truth and at accuracy, i.e. lower Brier scores, our credence functions ought to be probability functions, or so the norms of rationality demand.
This then answers the question whether synchronic logical and probabilistic coherence norms apply to language models, or at least whether they apply to them just as much as they apply to us. We argued that such norms do not always apply to pre-trained language models that simply model the distribution of words or tokens in a dataset of text. What is missing is the aim of being truthful. But this aim can be introduced in several ways, and with such an aim in place the rational norms do apply as a consequence. In particular, fine-tuned language models which are fine-tuned for truthfulness are under such coherence norms. In that case the norms of rationality apply to these models just as they apply to us. The next question to consider is whether this also holds for diachronic norms of rationality.

### v3

@@ 3 Coherence norms for language models
Belief in humans is subject to the norms of rationality. We are not always rational, but we ought to be, which is to say: we are under the norms of rationality, even if we don’t always live up to those requirements. Rationality itself is complex, including different branches or domains of rationality, each with its own set of norms. It is common to distinguish epistemic from practical rationality, with the former concerning what one ought to believe and the latter concerning what one ought to do. Our focus here is on epistemic rationality, but practical rationality might also be significant for language models, since such norms govern what one ought to do generally, including what speech acts one ought to perform in light of one’s beliefs and intentions. This form of rationality is widely discussed in speech act theory, [^6] See grice:studies and austin1962how for classic texts, and lorini:design for an application to language models. but not our topic here. In the following, we will not concentrate on what a model should do, assuming it has agency at all, nor on what it should say, assuming it makes assertions at all. Instead, we will focus on epistemic rationality: what the model should believe.
@@
The norms of epistemic rationality can themselves be analyzed as a series of distinct norms. One core part of these norms are coherence norms, which concern how the internal states of a system ought to relate to each other. Other norms of rationality concern the proper response to reasons or evidence. How the different parts of rationality relate to each other is itself controversial, but we will simply focus on coherence norms in the following. [^7] See kolodny:rational, kiesewetter:norm, worsnip:fitting and many others for the debate about whether coherence norms are derivative on norms tied to responding to reasons or whether they are an independent group of norms of rationality.
@@
Coherence norms come in different kinds, most notably as ones concerning full belief and ones concerning degrees of, or strength of belief. Most notably among the former are norms concerning logical coherence and among the latter are norms about probabilistic coherence. Logical coherence norms concern the relation of our beliefs to each other with respect to logical implication and consistency. To give a simple example, a logical coherence norm might demand that one ought not to believe a particular proposition $p$ as well as its negation $\neg p$. Probabilistic coherence norms in general also come in different kinds. One kind concerns belief about probabilities itself. For example, a norm might demand that one ought not to believe that the probability that it is going to rain is 90%, and also believe that the probability that it is not going to rain is 60%. But besides beliefs explicitly about probabilities, there are also coherence norms tied to subjective probabilities or credences. Credences correspond to how strongly one believes a particular proposition. So, to believe that Paris is in France with credence 99% is to believe it very strongly, but it is not itself a belief about probabilities. Therefore to believe that Paris is in France with 99% credence is not the same as to believe that the chance that Paris is in France is 99%. Credences should be coherent, and the view that they should satisfy the axioms of probability theory is generally called probabilism. For example, probabilism holds that if your credence that $p$ is $x$, then your credence that $\neg p$ ought to be (1x)(1-x), for $x\in[0,1]$. We will not discuss coherence norms tied to beliefs explicitly about probabilities in this paper, but focus on logical coherence and the coherence of credences instead.
@@
First, it is unclear why logical coherence should be required of language models. After all, they are machine learning models and thus they are supposed to learn the training data. But the training data for language models is generally incoherent. Language models are trained on text from the internet and other sources, which is not a coherent unified data set of text. Since the machine learning model is trained on incoherent text as its training data, why should we demand of it that it be coherent? In fact, coherence seems to go against what it is supposed to learn from the training data. Incoherent training data should lead to a model with incoherent beliefs if the model properly captures the training data. Thus to require the model to be coherent seems to be unjustified, and therefore the rational norm of coherence does not seem to apply to it. Unless the training data is coherent, the model should not be required to be coherent either. This contrasts with why coherence norms apply to human beings. We learn from a coherent world: a world without any contradictions. Thus our internal states should be coherent, to properly reflect what our minds are trained on. The world is mediated to us by evidence. But even if our evidence is conflicting, pointing partly in one direction and partly in another one, what the evidence points to — the world — is coherent. Our primary goal is not to learn the evidence, but to learn about the world from the evidence. That is why we need to resolve conflicting evidence. Thus the norm of logical coherence applies to us, since we aim to represent a coherent world. But no analogous case carries over directly to language models, whose primary goal is to learn the training data, not about the world.
@@ 3.1 Logical coherence norms
The solution to the first problem is similar to the argument given above in Section 2 for the coherence of beliefs in language models. It is true in general that a pretrained language model should not be seen as being under the norm of logical coherence. Such a model is simply aiming to model the training data, and that data is not logically coherent. To require coherence of a pretrained language model that is trained on incoherent data is to require something of it that goes against what it fundamentally aims to do: learn the training data. But things are again different for fine-tuned models, in particular ones that are fine-tuned for truthfulness. Properly fine-tuned models can acquire the function to produce truthful text, and if we assume that such models have internal representational states like beliefs and that those states are properly causally connected to the text they produce, then the aim of being truthful can support that logical coherence norms apply to the internal states of such models. Internal belief states that are incoherent will be contrary to the norm of truth which is part of the standard of correctness for belief itself. Thus inconsistent beliefs are in conflict with the standard of correctness of belief, and thus the coherence norm of consistency for beliefs applies. A pretrained model does not need to be under such norms, but a model fine-tuned for truthfulness is.
The norm of logical coherence is thus derivative on the norm of truthfulness. Any system that aims at truth must also aim at coherence, derivatively, since it cannot achieve its primary goal of truth without also being coherent. The truth norm is thus stronger than the coherence norm. Truth concerns the relationship between beliefs and the world. Coherence concerns the relationship between beliefs alone. The truth norm gives rise to the coherence norm since the world is coherent, and thus truth cannot be achieved without coherence. But coherence alone is not sufficient to characterize belief and what is distinctive of what kind of a state belief is. Beliefs that are perfectly coherent with each other but false are defective belief states: they are not how things should be and what is aimed for with these states being states of belief. A false belief is not how it should be, even if it is coherent with one’s other beliefs. The norm of truth is thus primary and the norm of logical coherence is derivative on it. Nonehteless, both norms apply, to use and to properly fine-tuned language models.
@@ 3.2 Defining credence for language models
which are assigned by the language model to the n-th token given that the previous tokens are $t_{1},...,t_{n-1}$, and turn that into an account of what the model’s credence in proposition $p$ is.
@@
We propose to solve this problem via the Minimal Assent Connection (MAC). The basic idea of MAC is simply to consider the probability of the model predicting the next token to be ‘yes’ when asked whether it is the case that $p$. This way next token probabilities can be directly tied to probabilities assigned to propositions. To put a label on it, we can say that the model minimally assents to $p$ just in case the probability it assigns to ‘yes’ being the next token after ‘Is it the case that $p$?’ is sufficiently high. This basic idea is promising, but to in order to turn it into a credence function that captures strength of belief more needs to be said and we will do this momentarily. But on the positive side, we can already note that this approach is completely general, since for any proposition $p$ we can consider the probabilities that the model assigns to the next token after the question whether it is the case that $p$. This probability comes directly from the language model. Next token probabilities are given in the output layer of the model and are thus easily accessible and require no special knowledge of the more hidden, internal layers of the model. This is just the kind of assignment of probability we want to exploit in assigning credences to the model. This assignment applies to any proposition $p$ uniformly. Furthermore, the probability of a ‘yes’ answer to the question whether $p$ is naturally associated with the strength of believing $p$. After all, it is natural to hold that the probability assigned to ‘yes’ after the question whether $p$ is closely correlated with the strength of believing $p$: the stronger the confidence of the model in $p$, the higher the probability for a ‘yes’ answer, all things being equal. But still, it can’t quite be a full answer and we can’t make this connection as outlined. That is, we can’t fully endorse that we can define $cr(p)$, the models credence in $p$, as follows:
cr(p)=P(Yes  Is it the case that p?)cr(p)=P($\text${Yes $\mid\text${ Is it the case that }p?) (1)
@@
There are essentially two flaws with this proposal. First, the model might assent to $p$ in ways other than with ‘yes’. It might also assent via other forms of affirmation, say ‘yeah’ or ‘indeed’. Using just (1) thus gives a credence that is too low. But we can simply fix this in one of two ways. First, we could force the model to answer either ‘yes’ or ‘no’, either by adding this requirement to the question prompt, or by constraining the model in other ways to do so (e.g. by constraining and re-normalizing the output distribution). However, adding such further constraints slightly changes the model or the text for which it predicts the next token. Making such a demand for either a ‘yes’ or a ‘no’ answer thus requires changes that make the method not completely general any more. It would not apply to any model any longer, only to ones properly modified to meet this constraint. But there is also a second way to solve this problem. We can leave the model and input alone, and instead simply sum over all affirmations, not just ‘yes’, but also ‘sure’, ‘yeah’, and so on, including multi-token sequences like ‘I couldn’t agree more.’ All such tokens or sequences of tokens we can label an assent sequence. If we call $AS$ the set of all such assent sequences, then we can better define the credence for $p$ by the model as the sum of all conditional probabilities of assenting. So, we can define the assent probability as:
@@
But we can fix the approach to overcome this difficulty. Strength of belief can be associated with how likely the model is to minimally assent when compared to how likely it is to minimally dissent, leaving aside other ways to respond to the question besides assenting or dissenting. So, we can simply consider the ratio of assenting to either assenting or dissenting. To make this idea work, we also need to consider the set of all dissent sequences $DS$, which are the ways the model might minimally dissent: ‘no’, ‘never’, etc..
@@
So understood, we should not expect that $as(p)$ and $ds(p)$ sum to 1, since the model might respond in other ways than assenting or dissenting. But we can simply renormalize by considering only the ratio of probabilities given to affirmations to those that are given to either affirmations or dissent, leaving out all other options. And this ratio is the proper way to assess the strength of belief of a model in a proposition $p$: how likely it is to assent when we consider only assent or dissent, leaving out all other options. This then is the proposal for how to assign credences to a language model in a way that is derivative only on the model’s own assignment of probabilities to the next token:
@@
We have not given a detailed list of assent and dissent sequences, and doing so requires one to make some choices that can be subject to debate. One concern is tied to context sensitivity. It could be that what counts as assent or dissent for a particular issue $p$ depends itself on $p$, and thus it is not clear whether assent and dissent sequences can simply be listed for questions concerning any $p$. Although this is in principle possible, it does not seem to be a major concern. First, it is unclear what plausible cases there are for this, and second it merely requires a reformulation of the proposal by relativizing assent and dissent sequences to a given $p$, and then consider the assent/dissent ratios concerning $p$ with each of those relativized to $p$ itself.
A more difficult case is presented by examples of sequences that concern the epistemic state of the model itself, as ‘It seems to me to be unlikely,’ ‘I am quite sure that $p$,’ or ‘it is doubtful.’ Such sequences contain ‘epistemic markers’ or ‘epistemic strengtheners or weakeners’ that report on the epistemic state of the model. Strong epistemic markers indicate that the model is in a strong epistemic position to make an assessment, as with ‘I am sure that…,’ or ‘I am quite certain that..’. Weak epistemic markers indicate a weak epistemic state, as in ‘I am not sure, but it seems that…’ or ‘I doubt that…’. Should sequences with such epistemic markers, either strong or weak, count as assent or dissent sequences for our purposes here? Should next token probabilities allocated to such sequences count towards assent or dissent probabilities?
@@
Whether or not present-day models’ credences match their epistemic markings is an empirical question. If they do they are epistemically better functioning than if they don’t. To determine whether there indeed is this correlation is subject to ongoing work. See also zhou-relying for more on models reporting on their own epistemic state.
Furthermore, we would like to note that
@@
Finally, we would like to make clear that the assignments of credences to beliefs for models as defined above are for revealed belief, since they are based on the model’s response to questions. This concerns belief as it is revealed by the model in its verbal behavior. Focusing on revealed belief does not mean that we take belief in general to be understood behavioristically. Behaviorist theories of belief and other mental states see those states as derivative on one’s behavior or dispositions to behave, rather than internal states that cause this behavior. [^8] Classic behaviorist accounts of belief include skinner1957verbal and ryle1949concept. The notion of revealed belief, that is, belief as it is revealed in one’s (verbal) behavior, makes sense whether or not belief in general is understood behavioristically. This approach is limited in that it does not work for models that have the ability to lie and thus answer questions contrary to what they believe. It is not clear how one can make sense of lying in language models, since it is not clear how to make sense of belief in language models besides revealed belief. To be clear, lying here is understood as not just saying something that is false but saying something that is contrary to what one believes. This might be a temporary limitation to understanding belief in language models. If so, then our account of credences for revealed belief can be understood as an account of evidence we get from verbal behavior for the credences of beliefs more generally. Verbal behavior provides defeasible evidence for what one believes, and our account would provide evidence for how strongly a model believes something. But until an account of model belief has been developed that goes beyond revealed belief, we take our account of credences to be an account of the strength of belief in language models.
@@ 3.3 The argument for probabilism
When focusing on synchronic coherence norms, this question quickly turns into the question whether the credence function should satisfying the axioms of probability theory and thus be a probability function. In the philosophical literature the view that a credence function should be a probability function is called probabilism. [^9] See joyce:prob, teitelbaum:bayes. Assuming credences for language models as spelled out above, does probabilism apply to them as well? Are language models under the coherence norm that their credence function satisfies the axioms of probability theory? For example, are they under the norm that $cr(p)=1-cr(\neg p)$?
To answer this question we first need to again distinguish pretrained language models from fine-tuned ones, in particular those fine-tuned for truthfulness. For a pretrained model we should not in general expect the model’s credences to obey the probability calculus, nor that they are under any norm to do so. Since the model will generally be trained on incoherent text, we can not expect that $cr(p)$, the credence assigned to $p$, is exactly $1-cr(\neg p)$. If the training data is sufficiently incoherent, then it can well be that the credence for each of them is over 50%. Since the model so far is merely a pure language model whose internal states are correct insofar they contribute to the training task and learning the data, we should not expect there to be a norm on these probability assignments that correspond to probabilism.
However, things are again different for the model that is fine-tuned on truthfulness. A model that aims at true beliefs is subject to the well-known accuracy arguments for probabilism. See joyce:prob, teitelbaum:bayes. The main idea of these arguments is as follows: if the goal is truth, then credences should be as close to the truth as possible: one’s credence in a true proposition should be close to 1, and one’s credence in a false proposition should be close to 0. However close one comes to this ideal can be seen as the accuracy of one’s credences. A standard measure of accuracy in this sense is the Brier score, basically the Euclidean distance between one’s credences in some propositions and their actual truth values (taken as 1 for true, and 0 for false) seen as points in a Euclidean space of the dimension that corresponds to the number of propositions, but without taking a square root. So, the Brier score for a credence function $cr$ from a finite set of propositions $p_{1},...p_{n}$ to [0,1] is defined as
@@
whereby $tv(p)=1$ if p is true and 0 if it is false. [^10] The Brier score is sometimes also given as the mean squared error, i.e. dividing the above by the dimension $n$. This difference does not matter for us here. The thought is that a credence function is better if its Brier score is lower, since one aims at the truth, and a lower Brier score of one’s credence function means that one’s credences are closer to how they should be. In joyce:prob, James Joyce showed that if one’s credence function does not meet the axioms of probability theory, then there is a different credence function that has a lower Brier score no matter what the world is like and which meets the axioms of probability theory. However, if one’s credence function does meet the axioms of probability theory, then this is not always the case. This argument is taken to support probabilism: A credence function that isn’t a probability function can always be dominated by one that is. Thus such a credence function isn’t what it should be, given that one aims at the truth. And thus probabilism is a rational norm that applies to one’s credences: one’s credence function ought to be a probability function. [^11] For further discussion of Joyce’s argument, see teitelbaum:bayes volume 2, pages 338-71. This argument applies to the credence functions of language models just as it applies to the credence functions of human beings. As long as we aim at the truth and at accuracy, i.e. lower Brier scores, our credence functions ought to be probability functions, or so the norms of rationality demand.
This then answers the question whether synchronic logical and probabilistic coherence norms apply to language models, or at least whether they apply to them just as much as they apply to us. We argued that such norms do not always apply to pretrained language models that simply model the distribution of words or tokens in a dataset of text. What is missing is the aim of being truthful. But this aim can be introduced in several ways, and with such an aim in place, the rational norms do apply as a consequence. In particular, fine-tuned language models which are fine-tuned for truthfulness are under such coherence norms. In that case, the norms of rationality apply to these models just as they apply to us. The next question to consider is whether this also holds for diachronic norms of rationality.